Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 194     Entry time: 08 May 2023, 02:38
Author: Jan-Etienne Pudell 
Group: MID 
Subject: Shift summary 

Stable beam delivery with 23keV and 140µJ at 2.2, 1.1 and 0.55MHz

 

User data acquisition during the day interrupted by storage and detector DAQ problems:

- problem with detector DAQ, which is correlated to the full SSD cache.

- we run out of storage, this problem was solved by the DAQ problems. We lost 6h of measurement time today, due to storage limitations. 

- 2 detector modules stopped sending data

ELOG V3.1.4-7c3fd00